How would I solve for x in the problem 6x-3y =54
Found 2 solutions by jorel555, addingup:
Answer by jorel555(1290)

6x-3y=54
6x=54+3y
x=9+1/2y is your result when solving for x!!!!!!!!!!

Answer by addingup(3677)

6x-3y = 54 divide both sides by 6
x-3/6y = 9
x = 9+3/6y